OpenStudy (anonymous):

What is the center of the ellipse defined by the equation (x+8)^2/36 + (y-7)^2/9=1?

OpenStudy (anonymous):

Put x+8 = X y-7 = Y so the New equation becomes X^2 / 36 + Y^2/9 = 1 with Centre (0,0) ----> X=0 Y=0 so x+8 = X y-7 = Y x = -8 y = 7 So centre (-8,7)